Assume that each iron atom has a permanent magnetic moment equal to 2 Bohr magnetons (1 Bohr magnetic equals 9.27 × 10^-24 A-m^2). The density of atoms in iron is 8.52 × 10^28 atoms/m^3. (a) Find the maximum magnetization I in a long cylinder of iron. (b) Find the maximum magnetic field B on the axis inside the cylinder.

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. f = 8.52 x 10 ^28 atoms / m^3 For maximum ‘I’, Let us consider the no. of atoms present in 1 m^3 of volume. Given: m per atom = 2 × 9.27 × 10^–24 A–m^2 I = net m / V = 2 × 9.27 × 10^–24 × 8.52 × 10^28 ≈ 1.58 × 10^6 A/m B = μ0 (H + I) = μ0 [∴ H = 0 in this case] = 4π × 10^–7 × 1.58 × 10^6 = 1.98 × 10^–1 ≈ 2.0 T